Reports 1-1 of 1 Clear search Modify search
Injection system (Laser power stabilization)
derossi, chiummo - 17:33 Tuesday 13 August 2019 (46660) Print this report
PSTAB noise projection

Today we performed a laser intensity noise projection on Hrec by opening the PSTAB loop at 1249742972 and closing it at 1249743115. We took as a calm period for the projection the gps 1249742750, during 60 s.

Compared to the last projection (46453) the level of PD1 (out of loop error signal) is very different: it is now a factor ~200 below the sensitivity curve. This might be explained by a different working point of the ITF (to be checked), since the last projection was done during the single B1 PD operation.

Images attached to this report
Comments to this report:
derossi - 10:00 Wednesday 14 August 2019 (46664) Print this report

In the projection reported in 46453 there was an error on the frequency axis, which was shifted by a factor 2. I attach here the corrected plot.

Images attached to this comment
Search Help
×

Warning

×